If 2x-3(x+4)=-5,then x=
2x-3x-12=-5 (use distributive property)
-x-12+12=-5+12 (like terms)
-x=7
x=-7

Answer by aves(2)

2x-3x-12=-5 (use distributive property)
-x-12+12=-5+12 (like terms)
-x=7
x=-7